OpenStudy (anonymous):

Prove the following identities: 1.) (tan^2 x) (sin^2 x) + sin^2 x = tan^2 x 2.) sin^2 x + cos^2 x = csc^2 x + sin^2 x 3.) (sin x + cos x) sin x = 1 - cos x (cos x - sin x) HELP ME PLEASE.. :D

hartnn (hartnn):

factor out sin^2x from the first equation's left side, what you get ?

OpenStudy (anonymous):

how to factor po?

hartnn (hartnn):

(tan^2 x) (sin^2 x) + sin^2 x =[(tan^2 x)+1] (sin^2 x) =[.......] sin^2 x can you fill the blank using a pythagorean identity ?

OpenStudy (anonymous):

hm, yes po. the answer is sec^2 x ?

hartnn (hartnn):

yes, so sec^2x * sin^2 x =... ?

OpenStudy (anonymous):

hm, i don't know the answer. hm, is sec^2x * sin^2 x has a formula?

hartnn (hartnn):

no, but sec^2x = 1/cos^2 x and sin^2x /cos^2x = tan^2 x which proves the first identity, got it ?

OpenStudy (anonymous):

no .. i didn't get. im so sorry. im so slow. ahm , how did you get [(tan^2 x)+1] (sin^2 x)

hartnn (hartnn):

by factoring out sin^2x its similar to (s+st) = s(1+t)

OpenStudy (anonymous):

ahh. :D now i know, i get it .. but how abou this : sec^2x * sin^2 x ?

hartnn (hartnn):

sec = 1/cos so, sec^2x * sin^2 = sin^2 / cos^2 and tan=sin / cos so, sin^2 / cos^2 = tan^2 = right side.

OpenStudy (anonymous):

ahh.. i get it .. :D thank you po . :D hm, how about #2? will you help me on that?

hartnn (hartnn):

#2 isn't an identity, it can't be proved. unless you have typed the question incorrectly...

hartnn (hartnn):

i can help you with #3 , if you still need it.

OpenStudy (anonymous):

ahh .. that's why i can't prove it .. thank you po . hm, yes .. #3 pls.

hartnn (hartnn):

do you know, sin^2 +cos^2 =.... ?

OpenStudy (anonymous):

yes .. it is equal to 1.

hartnn (hartnn):

good, so from that sin^2+cos^2=1 what will sin^2 x =... ?

OpenStudy (anonymous):

ahm , 1 - cos^2x ? am i right?

hartnn (hartnn):

correct :)

hartnn (hartnn):

(sin x + cos x) sin x = sin^2x + sin x cos x now put here, the value of sin^2x you got,

OpenStudy (anonymous):

1 - cos^2x + sin x cos x = 1 - cos x (cos x - sin x)

hartnn (hartnn):

good! so you got 3rd by yourself :) nice work...

OpenStudy (anonymous):

ahh .. is that all only ?

hartnn (hartnn):

es, we proved left side = (sin x + cos x) sin x = sin^2x + sin x cos x =1 - cos^2x + sin x cos x = 1 - cos x (cos x - sin x) =right side so its an identity.....

hartnn (hartnn):

and bdw, i see this is your first day here, so \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

OpenStudy (anonymous):

ahh.. thank you so much po. :D hm, it's not my first day here po , i just make a new account po because i can't use the other one . :D but, thank you po for noticed me .. :D

hartnn (hartnn):

can i ask you a question ? whats a 'po' ?

OpenStudy (anonymous):

ahm sure po . hehe .. i'm a FILIPINO po then we used "po" po in respecting someone ..

hartnn (hartnn):

ohhh..thanks :) can i ask you one more question ? is you real name Quennie ? if you don't want to, don't answer...

hartnn (hartnn):

aaah...i see you have updated your profile, and it is :)

OpenStudy (anonymous):

hm, always welcome po. hehe, yes. that's my real name .. im QUENNIE DIMAANO po. haha! :D thank you again po.

hartnn (hartnn):

just so that you know, that i am also a math lover :)

OpenStudy (anonymous):

ahh .. yes . . i already know that po .. :D